# ConceptDrift: Uncovering Biases through the Lens of Foundation Models

### v2

@@ Appendix A Zero-Shot Prompts
In Tab. A.1, we structured the class names used for initializing the initial weights of the linear layer, along with the prompt templates employed in the zero-shot classification experiments discussed in Sec. 5.4.
Table A.1: Class names and prompts used in the zero-shot classification task.
@@
bias-enhanced prompt a photo of a {cls} in the {bias} a photo of a {bias} with {cls} a/an {cls} comment about {bias}
Appendix B Biases and top class-neutral concepts
@@
We present the exhaustive list of biases for the Waterbirds (Tab. B.1 & B.2), CelebA (Tab. B.3 & B.4) and CivilComments (Tab. B.5 & B.6) datasets. We also present top class-neutral concepts for ImageNet classes. Concept filtering on ImageNet was performed using WordNet relationships alone, without the intervention of a Large Language Model. Accounting for the size of the dataset, we will publish the available data on our repository upon acceptance, and we will restrict the presentation within the context of the current format to a few classes for illustrative purposes in Tab. B.7 -B.15.
Table B.1: Top Waterbirds class-neutral concepts for ”landbird”.
@@
Table B.2: Top Waterbirds class-neutral concepts for ”waterbird”.
@@
Table B.3: Top CelebA class-neutral concepts for ”non-blonde”.
@@
Table B.4: Top CelebA class-neutral concepts for ”blonde”.
Blonde Score
yellow -0.021363914
white -0.04120779
golden -0.041870475
woman with blowing -0.042802572
woman with laughing -0.046326935
girl with blue eyes -0.050636113
woman with playing -0.0510059
girl next door -0.05101496
woman with green -0.05205536
sweater -0.054968417
gold -0.057154775
scarf -0.0585289
pale -0.05990547
woman with blue eyes -0.05999875
tank top -0.060728252
Table B.5: Top CivilComments class-neutral concepts for ”non-offensive”.
@@
Table B.6: Top CivilComments class-neutral concepts for ”offensive”.
@@
die on the golf -0.001492977
Muslim Canadians Party -0.002062559
serve them at the lunch -0.00358063
culture is no worse -0.005233765
banned -0.006861687
pretend they are all goodness -0.007072449
pop that particular rainbow -0.008466244
member becomes so sick -0.008648455
men and she wanted -0.008733362
Canadian Muslims -0.009060442
destroy -0.010027111
six Baltic -0.01069364
customers when it not acceptable -0.011366904
Table B.7: Top ImageNet class-neutral concepts for ”American Lobster”.
”American Lobster” – Top Concepts Score
pasta with shrimp 0.3148532
adirondack sitting 0.31178916
large shrimp 0.3085378
shrimp and pasta 0.30821544
shrimp cooking 0.3032636
large cast cooking 0.30282205
pasta with shrimp and cheese 0.2984723
adirondack 0.29169592
and mussels 0.2904386
legs and other seafood 0.28870153
close up of a shrimp 0.2871693
of pasta with shrimp 0.2823377
seafood 0.2808096
mussels 0.28018713
cast cooking 0.27951774
shrimp and cheese 0.27841187
shrimp 0.2726224
Table B.8: Top ImageNet class-neutral concepts for ”Bald Eagle”.
”Bald Eagle” – Top Concepts Score
osprey flying 0.2888234
osprey 0.27377927
row of american flags 0.26513425
group of american flags flying 0.25681192
american flag and american flag 0.23945728
emu standing 0.23752406
yellowstone national 0.23740456
Table B.9: Top ImageNet class-neutral concepts for ”Ballpoint Pen”.
”Ballpoint Pen” – Top Concepts Score
wearing a pilot 0.36252645
markers 0.36050144
sharpie 0.34563553
stylus 0.34399948
pair of eyeglasses 0.32426757
notepad 0.3228797
calligraphy 0.3222967
paint and markers 0.32198787
close up of a needle 0.32131955
on a straw 0.3209229
and markers 0.32090995
eyeglasses 0.31017447
dots 0.30764964
crayons 0.30598855
Table B.10: Top ImageNet class-neutral concepts for ”Crossword”.
@@
Table B.11: Top ImageNet class-neutral concepts for ”Coffeemaker”.
@@
Table B.12: Top ImageNet class-neutral concepts for ”Doormat”.
@@
Table B.13: Top ImageNet class-neutral concepts for ”Eraser”.
@@
Table B.14: Top ImageNet class-neutral concepts for ”Fire Truck”.
@@
Table B.15: Top ImageNet class-neutral concepts for ”Guacamole”.
”Guacamole” – Top Concepts Score
tomatoes and avocado 0.33948907
avocado 0.3125001
nachos 0.30761188
ham and parsley 0.2960047
with avocado 0.2952027
colorful mexican 0.28313732
bacon and parsley 0.27773544
of avocado 0.2754345
side of salsa 0.27470407
peridot 0.2734925
salsa 0.27124816
nachos with 0.27092364
lime body 0.27088284
cheese and parsley 0.26951405
pile of limes 0.26944226
peas and bacon 0.2671204
limes and limes 0.26682717
lime cut 0.2662533
nachos with and 0.26580203
pasta with peas 0.2626204
tacos 0.2619322
pasta with ham and parsley 0.26183394
tomatoes and cilantro 0.26161948
Appendix C Algorithm
We present the formal implementation of our proposed ConceptDrift approach in Alg.  1. We annotate the main steps presented in Sec. 3. At line 2, we initialize the class weights of our linear probing layer with the text embeddings of the class names, i.e. the zero-shot classification weights of the foundation model $M$. We then fine-tune the layer, at line 3, on the given dataset. At lines 5 and 6, we filter the list of concepts and compute the embeddings for the remaining class-neutral concepts. The filtering can be performed by means of employing WordNet associations, Large Language Models, or both. With the embeddings of the class-neutral concepts at hand, we proceed to determine the biases for each class. At line 7, we initialize our set of biases with an empty set, and we compute the similarities between each selected concept and each class at line 8.
We further proceed to iterate through each class (line 9). For a given class $k$, we compute the set of scores for all concepts with respect to class $k$ and store it in $s_{k}$, at line 11. For each class-neutral concept, its ranking score for the current class is the difference between its similarity to the current class and the smallest similarity with a different class. Lines 12-14 formally implement the dynamic thresholding procedure proposed in Sec. 3. Finally, we select the concepts with ranking scores above the computed threshold and store them as biases for class $k$, at lines 15 and 16.
Algorithm 1 ConceptDrift

### v3

@@ Appendix A Zero-Shot Prompts
In Tab. 7, we structured the class names used for initializing the initial weights of the linear layer, along with the prompt templates employed in the zero-shot classification experiments discussed in Sec. 5.3.
Table 7: Class names and prompts used in the zero-shot classification task.
@@
SC-enhanced prompt a photo of a {cls} in the {SC} a photo of a {SC} with {cls} a/an {cls} comment about {SC}
Appendix B Results of state-of-the-art models
We provide results using the same data and experimental setup used for Tables 1 and 2, for an exhaustive list of ImageNet classifiers, in Tables 8, 9 and 10. Pre-trained models together with their respective weight sets are employed from the torchvision package.
Table 8: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab. 1. As with Fig. 1 and Fig. 4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
Prompt employed (correct class highlighted in bold, SC in red)
Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
alexnet - V1 (Krizhevsky, 2014) 100.0 4.6 96.2 23.3
convnext_tiny - V1 (Liu et al., 2022) 100.0 77.2 94.2 69.8
convnext_small - V1 (Liu et al., 2022) 100.0 92.9 96.3 78.3
convnext_base - V1 (Liu et al., 2022) 100.0 83.5 99.3 78.0
convnext_large - V1 (Liu et al., 2022) 100.0 88.2 99.6 81.9
densenet121 - V1 (Huang et al., 2016) 100.0 52.3 93.1 74.8
densenet161 - V1 (Huang et al., 2016) 100.0 52.8 84.0 51.6
densenet201 - V1 (Huang et al., 2016) 100.0 49.7 86.1 80.8
efficientnet_b0 - V1 (Tan & Le, 2019) 100.0 64.5 99.2 92.9
efficientnet_b1 - V1 (Tan & Le, 2019) 100.0 42.6 88.1 67.1
efficientnet_b1 - V2 (Tan & Le, 2019) 100.0 84.2 99.9 69.5
efficientnet_b2 - V1 (Tan & Le, 2019) 100.0 61.7 99.6 82.4
efficientnet_b3 - V1 (Tan & Le, 2019) 100.0 89.1 99.1 92.8
efficientnet_b4 - V1 (Tan & Le, 2019) 100.0 94.4 99.9 72.7
efficientnet_b5 - V1 (Tan & Le, 2019) 100.0 82.9 99.4 86.3
efficientnet_b6 - V1 (Tan & Le, 2019) 100.0 91.2 100.0 95.6
efficientnet_b7 - V1 (Tan & Le, 2019) 100.0 88.3 99.9 90.3
efficientnet_v2_s - V1 (Tan & Le, 2021) 100.0 98.3 99.2 88.1
efficientnet_v2_m - V1 (Tan & Le, 2021) 100.0 95.1 99.8 94.3
efficientnet_v2_l - V1 (Tan & Le, 2021) 100.0 96.0 99.1 83.6
googlenet - V1 (Szegedy et al., 2015) 100.0 45.4 95.2 57.0
inception_v3 - V1 (Szegedy et al., 2016) 100.0 82.2 98.9 80.8
maxvit_t - V1 (Tu et al., 2022) 100.0 91.7 99.7 85.3
mnasnet0_5 - V1 (Tan et al., 2019) 100.0 23.3 96.7 60.8
mnasnet0_75 - V1 (Tan et al., 2019) 100.0 36.9 98.4 71.1
mnasnet1_0 - V1 (Tan et al., 2019) 100.0 32.0 89.4 74.9
mnasnet1_3 - V1 (Tan et al., 2019) 100.0 63.9 91.7 75.6
mobilenet_v2 - V1 (Sandler et al., 2018) 100.0 26.2 84.0 45.0
mobilenet_v2 - V2 (Sandler et al., 2018) 100.0 54.1 98.6 62.8
mobilenet_v3_small - V1 (Howard et al., 2019) 100.0 16.6 91.0 30.3
mobilenet_v3_large - V1 (Howard et al., 2019) 100.0 25.4 94.2 23.1
mobilenet_v3_large - V2 (Howard et al., 2019) 100.0 56.0 96.6 73.7
Table 9: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab. 1. As with Fig. 1 and Fig. 4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
Prompt employed (correct class highlighted in bold, SC in red)
Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
regnet_y_400mf - V1 (Radosavovic et al., 2020) 100.0 35.6 74.7 34.7
regnet_y_400mf - V2 (Radosavovic et al., 2020) 100.0 72.1 97.7 87.9
regnet_y_800mf - V1 (Radosavovic et al., 2020) 100.0 22.7 94.8 54.4
regnet_y_800mf - V2 (Radosavovic et al., 2020) 100.0 81.7 98.9 88.0
regnet_y_1_6gf - V1 (Radosavovic et al., 2020) 100.0 47.0 96.4 71.7
regnet_y_1_6gf - V2 (Radosavovic et al., 2020) 100.0 88.2 96.4 74.1
regnet_y_3_2gf - V1 (Radosavovic et al., 2020) 100.0 33.5 99.4 90.9
regnet_y_3_2gf - V2 (Radosavovic et al., 2020) 100.0 94.6 98.4 75.4
regnet_y_8gf - V1 (Radosavovic et al., 2020) 100.0 58.3 71.3 54.8
regnet_y_8gf - V2 (Radosavovic et al., 2020) 100.0 98.2 96.8 67.5
regnet_y_16gf - V1 (Radosavovic et al., 2020) 100.0 86.7 97.8 49.8
regnet_y_16gf - V2 (Radosavovic et al., 2020) 100.0 98.7 91.8 83.5
regnet_y_16gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 97.7 78.2
regnet_y_16gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 78.3 100.0 90.7
regnet_y_32gf - V1 (Radosavovic et al., 2020) 100.0 84.5 99.1 82.3
regnet_y_32gf - V2 (Radosavovic et al., 2020) 100.0 98.0 99.5 79.6
regnet_y_32gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 93.5 71.5
regnet_y_32gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 97.2 100.0 85.6
regnet_y_128gf - SWAG_E2E_V1 (Radosavovic et al., 2020) 100.0 99.6 54.4 67.8
regnet_y_128gf - SWAG_LINEAR_V1 (Radosavovic et al., 2020) 100.0 90.8 99.9 96.9
regnet_x_400mf - V1 (Radosavovic et al., 2020) 100.0 37.4 82.3 29.0
regnet_x_400mf - V2 (Radosavovic et al., 2020) 100.0 50.7 99.5 81.7
regnet_x_800mf - V1 (Radosavovic et al., 2020) 100.0 25.4 77.0 52.6
regnet_x_800mf - V2 (Radosavovic et al., 2020) 100.0 75.5 97.1 71.4
regnet_x_1_6gf - V1 (Radosavovic et al., 2020) 100.0 38.2 76.9 66.7
regnet_x_1_6gf - V2 (Radosavovic et al., 2020) 100.0 82.2 99.2 88.1
regnet_x_3_2gf - V1 (Radosavovic et al., 2020) 100.0 45.1 62.1 69.4
regnet_x_3_2gf - V2 (Radosavovic et al., 2020) 100.0 83.4 99.6 88.3
regnet_x_8gf - V1 (Radosavovic et al., 2020) 100.0 41.8 98.8 89.0
regnet_x_8gf - V2 (Radosavovic et al., 2020) 100.0 93.5 99.2 81.6
regnet_x_16gf - V1 (Radosavovic et al., 2020) 100.0 48.7 86.6 54.5
regnet_x_16gf - V2 (Radosavovic et al., 2020) 100.0 93.4 97.9 88.1
regnet_x_32gf - V1 (Radosavovic et al., 2020) 100.0 66.1 85.9 46.0
regnet_x_32gf - V2 (Radosavovic et al., 2020) 100.0 97.4 99.6 83.7
resnet18 - V1 (He et al., 2016) 100.0 36.8 84.3 41.9
resnet34 - V1 (He et al., 2016) 100.0 32.9 54.1 26.1
resnet50 - V1 (He et al., 2016) 100.0 30.1 73.9 54.5
resnet50 - V2 (He et al., 2016) 100.0 80.1 99.7 88.4
resnet101 - V1 (He et al., 2016) 100.0 60.4 92.3 84.4
resnet101 - V2 (He et al., 2016) 100.0 93.4 98.8 87.4
resnet152 - V1 (He et al., 2016) 100.0 66.1 98.4 78.2
resnet152 - V2 (He et al., 2016) 100.0 93.1 98.5 92.5
resnext50_32x4d - V1 (Xie et al., 2017) 100.0 45.5 92.6 74.8
resnext50_32x4d - V2 (Xie et al., 2017) 100.0 80.3 98.5 88.0
resnext101_32x8d - V1 (Xie et al., 2017) 100.0 66.6 84.7 61.2
resnext101_32x8d - V2 (Xie et al., 2017) 100.0 90.3 99.4 85.2
resnext101_64x4d - V1 (Xie et al., 2017) 100.0 77.9 97.7 74.2
Table 10: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab. 1. As with Fig. 1 and Fig. 4, we note that the performance of these models is significantly affected, even though the correct class is clearly illustrated right in front and center while, and the predicted class is absent from the generated images.
Prompt employed (correct class highlighted in bold, SC in red)
Model - Weights a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
shufflenet_v2_x0_5 - V1 (Ma et al., 2018) 100.0 23.8 36.9 21.0
shufflenet_v2_x1_0 - V1 (Ma et al., 2018) 99.8 30.4 72.1 55.9
shufflenet_v2_x1_5 - V1 (Ma et al., 2018) 100.0 41.2 97.9 52.5
shufflenet_v2_x2_0 - V1 (Ma et al., 2018) 100.0 61.4 99.3 64.5
squeezenet1_0 - V1 (Iandola et al., 2016) 100.0 11.4 95.9 28.7
squeezenet1_1 - V1 (Iandola et al., 2016) 100.0 13.8 91.2 46.1
swin_t - V1 (Liu et al., 2021b) 100.0 72.7 96.9 81.6
swin_s - V1 (Liu et al., 2021b) 100.0 74.3 99.3 81.4
swin_b - V1 (Liu et al., 2021b) 100.0 81.5 95.2 72.6
swin_v2_t - V1 (Liu et al., 2021b) 100.0 76.2 88.7 73.6
swin_v2_s - V1 (Liu et al., 2021b) 100.0 85.7 90.7 74.4
swin_v2_b - V1 (Liu et al., 2021b) 100.0 73.0 96.0 85.2
vgg11 - V1 (Simonyan & Zisserman, 2015) 100.0 9.9 96.6 60.7
vgg11_bn - V1 (Simonyan & Zisserman, 2015) 100.0 15.9 86.7 54.8
vgg13 - V1 (Simonyan & Zisserman, 2015) 100.0 5.9 97.7 68.1
vgg13_bn - V1 (Simonyan & Zisserman, 2015) 100.0 15.5 87.0 13.0
vgg16 - V1 (Simonyan & Zisserman, 2015) 100.0 12.2 93.7 66.0
vgg16_bn - V1 (Simonyan & Zisserman, 2015) 100.0 22.6 96.1 70.7
vgg19 - V1 (Simonyan & Zisserman, 2015) 100.0 26.6 98.5 40.8
vgg19_bn - V1 (Simonyan & Zisserman, 2015) 100.0 35.9 83.1 46.2
vit_b_16 - V1 (Dosovitskiy et al., 2021) 100.0 85.4 97.1 75.2
vit_b_16 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 88.9 95.8 59.1
vit_b_16 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 79.2 99.9 84.6
vit_b_32 - V1 (Dosovitskiy et al., 2021) 100.0 56.1 96.0 86.0
vit_l_16 - V1 (Dosovitskiy et al., 2021) 100.0 55.9 95.3 76.0
vit_l_16 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 92.1 100.0 93.3
vit_l_16 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 97.4 100.0 72.5
vit_l_32 - V1 (Dosovitskiy et al., 2021) 100.0 54.0 96.9 82.9
vit_h_14 - SWAG_E2E_V1 (Dosovitskiy et al., 2021) 100.0 99.4 98.5 86.4
vit_h_14 - SWAG_LINEAR_V1 (Dosovitskiy et al., 2021) 100.0 99.7 100.0 96.1
wide_resnet50_2 - V1 (Zagoruyko & Komodakis, 2016) 100.0 60.6 95.7 63.9
wide_resnet50_2 - V2 (Zagoruyko & Komodakis, 2016) 100.0 83.9 99.5 85.4
wide_resnet101_2 - V1 (Zagoruyko & Komodakis, 2016) 100.0 69.3 84.1 72.8
wide_resnet101_2 - V2 (Zagoruyko & Komodakis, 2016) 100.0 91.0 98.8 84.9
Appendix C SCs and top class-neutral concepts
@@
We present the exhaustive list of SCs found for the Waterbirds (Tab. 12 & 12), CelebA (Tab. 13) and CivilComments (Tab. 15 & 15) datasets. We also present top class-neutral concepts for ImageNet classes. Concept filtering on ImageNet was performed using WordNet relationships alone, without the intervention of a Large Language Model. Accounting for the size of the dataset, we will publish the available data on our repository upon acceptance, and we will restrict the presentation within the context of the current format to a few classes for illustrative purposes in Tab. 20 -19.
Table 11: Top Waterbirds class-neutral concepts for ”landbird”.
@@
Table 12: Top Waterbirds class-neutral concepts for ”waterbird”.
@@
Table 13: Top CelebA class-neutral concepts for ”non-blonde”.
@@
Table 14: Top CivilComments class-neutral concepts for ”non-offensive”.
@@
Table 15: Top CivilComments class-neutral concepts for ”offensive”.
@@
Table 16: Top ImageNet class-neutral concepts for ”Crossword”.
@@
Table 17: Top ImageNet class-neutral concepts for ”Guacamole”.
”Guacamole” – Top Concepts Score
tomatoes and avocado 0.33948907
avocado 0.3125001
nachos 0.30761188
ham and parsley 0.2960047
with avocado 0.2952027
colorful mexican 0.28313732
bacon and parsley 0.27773544
of avocado 0.2754345
side of salsa 0.27470407
peridot 0.2734925
salsa 0.27124816
nachos with 0.27092364
lime body 0.27088284
cheese and parsley 0.26951405
pile of limes 0.26944226
peas and bacon 0.2671204
limes and limes 0.26682717
lime cut 0.2662533
nachos with and 0.26580203
pasta with peas 0.2626204
tacos 0.2619322
pasta with ham and parsley 0.26183394
tomatoes and cilantro 0.26161948
Table 18: Top ImageNet class-neutral concepts for ”Bald Eagle”.
”Bald Eagle” – Top Concepts Score
osprey flying 0.2888234
osprey 0.27377927
row of american flags 0.26513425
group of american flags flying 0.25681192
american flag and american flag 0.23945728
emu standing 0.23752406
yellowstone national 0.23740456
Table 19: Top ImageNet class-neutral concepts for ”Ballpoint Pen”.
”Ballpoint Pen” – Top Concepts Score
wearing a pilot 0.36252645
markers 0.36050144
sharpie 0.34563553
stylus 0.34399948
pair of eyeglasses 0.32426757
notepad 0.3228797
calligraphy 0.3222967
paint and markers 0.32198787
close up of a needle 0.32131955
on a straw 0.3209229
and markers 0.32090995
eyeglasses 0.31017447
dots 0.30764964
crayons 0.30598855
Table 20: Top ImageNet class-neutral concepts for ”American Lobster”.
”American Lobster” – Top Concepts Score
pasta with shrimp 0.3148532
adirondack sitting 0.31178916
large shrimp 0.3085378
shrimp and pasta 0.30821544
shrimp cooking 0.3032636
large cast cooking 0.30282205
pasta with shrimp and cheese 0.2984723
adirondack 0.29169592
and mussels 0.2904386
legs and other seafood 0.28870153
close up of a shrimp 0.2871693
of pasta with shrimp 0.2823377
seafood 0.2808096
mussels 0.28018713
cast cooking 0.27951774
shrimp and cheese 0.27841187
shrimp 0.2726224
Table 21: Top ImageNet class-neutral concepts for ”Coffeemaker”.
@@
Table 22: Top ImageNet class-neutral concepts for ”Doormat”.
@@
Table 23: Top ImageNet class-neutral concepts for ”Eraser”.
@@
Table 24: Top ImageNet class-neutral concepts for ”Fire Truck”.
@@
Appendix D Algorithm
We present the formal implementation of our proposed WASP approach in Alg.  1. We annotate the main steps presented in Sec. 3. At line 2, we initialize the class weights of our linear probing layer with the text embeddings of the class names, i.e. the zero-shot classification weights of the foundation model $M$. We then fine-tune the layer, at line 3, on the given dataset. At lines 5 and 6, we filter the list of concepts and compute the embeddings for the remaining class-neutral concepts. The filtering can be performed by means of employing WordNet associations, Large Language Models, or both. With the embeddings of the class-neutral concepts at hand, we proceed to determine the SCs for each class. At line 7, we initialize our set of spuriously correlated concepts with an empty set, and we compute the similarities between each selected concept and each class at line 8.
We further proceed to iterate through each class (line 9). For a given class $k$, we compute the set of scores for all concepts with respect to class $k$ and store it in $s_{k}$, at line 11. For each class-neutral concept, its ranking score for the current class is the difference between its similarity to the current class and the smallest similarity with a different class. Lines 12-14 formally implement the dynamic thresholding procedure proposed in Sec. 3. Finally, we select the concepts with ranking scores above the computed threshold and store them as spuriously correlated concepts for class $k$, at lines 15 and 16.
Algorithm 1 WASP